OpenStudy (ashtorah13):

Last one, What is the cost in dollars to steam clean a room W yards wide and L yards long it the steam cleaners charge 10 cents per square foot? A. 0.9WL B. 0.3WL C. 0.1WL D. 9WL E. 3WL

OpenStudy (ashtorah13):

my guess is D

OpenStudy (photon336):

now we can imagine that we have a box right? and the total area will be L*W square yards. now if we know that we have 10 cents per square foot right? that would be 10 cents per (1/9) square yards. the way to represent the total cost would be L*W \[(\frac{ 0.10~cents }{ \frac{ 1 }{ 9 } yard^{2} })*L*W (yard^{2}) = cost \] \[(\frac{ 0.10~cents }{ \frac{ 1 }{ 9 } \cancel\yard^{2} })*L*W (\cancel\yard^{2}) = cost \]

OpenStudy (ashtorah13):

So a?

OpenStudy (photon336):

but do you see how the units cancel out?

OpenStudy (photon336):

@Ashtorah13 this is probably one of the most important parts to problem solving

OpenStudy (ashtorah13):

Ya I do

OpenStudy (photon336):

yards cancel out. and we're left with the cost in cents. another important thing I did was convert square feet to square yards.

OpenStudy (photon336):

\[\frac{ \frac{ 1 }{ 10 } }{ \frac{ 1 }{ 9 } } = \frac{ 9 }{ 10 }\]

OpenStudy (photon336):

If you're interested in the whole complex fraction \[\frac{ \frac{ 1 }{ 10 }*90 }{ \frac{ 1 }{ 9 }*90 } = \frac{ \frac{ 90 }{ 10 } }{ 10 } = \frac{ 9 }{ 10 }\]

OpenStudy (ashtorah13):

HAHA I seem to be lacking my basic math tonight. I had to rethink how you got 9/10 then figured it out cause when you divide you flip the second fraction and multiply them straight across.

OpenStudy (photon336):

check out how I did it. you can use a complex fraction or just do this @Ashtorah13

OpenStudy (photon336):

\[\frac{ 1 }{10 } \div \frac{ 9 }{ 10 }\] KEEP THE FIRST FRACTION CHANGE THE SIGN TO MULTIPLICATION FLIP THE LAST FRACTION \[\frac{ 1 }{ 10 }*(\frac{ 1 }{ 9}) = \frac{ 1 }{ 10 }*(\frac{ 9 }{ 1 }) = \frac{ 9 }{ 10 }\]

OpenStudy (ashtorah13):

I liked the less complex one. but since 9/10 is .9 it would be answer A cause we still cant sold lw so it would be .9LW

OpenStudy (photon336):

Sorry that should be (1/9)

OpenStudy (photon336):

yep exactly, 0.9LW

OpenStudy (ashtorah13):

Thanks so much!

OpenStudy (photon336):

yep the second method i used was actually easier and less work

OpenStudy (ashtorah13):

I am the type that needs the work for problems to make since.

OpenStudy (photon336):

if you're stuck try this one

OpenStudy (photon336):

\[\frac{ 1 }{ 10 } \div \frac{ 1 }{ 3 }\]

OpenStudy (ashtorah13):

3/10

OpenStudy (ashtorah13):

or .3

OpenStudy (photon336):

how did you get that?

OpenStudy (ashtorah13):

you flip the second fraction them multiply. Thats how i remember "dividing" fractions was taught to me in high school anyway.

OpenStudy (photon336):

always good to refresh on this stuff

OpenStudy (photon336):

it seems trivial but it will help you greatly

OpenStudy (ashtorah13):

Definitely. Oh no I like it. I am trying to get into collage.

OpenStudy (photon336):

another thing is to solve equations first before plugging in the numbers like identify the variable you want to solve for FIRST.

OpenStudy (photon336):

look into dimensional analysis, algebra I.

OpenStudy (ashtorah13):

will do. I am off to bed to absorb all i have relearned. Thanks again!

OpenStudy (photon336):

np take care